Given that a-3b/a+2b=3, find the value of a+3b/a-2b. Express your answer as an integer or common or improper fraction reduced to lowest terms. Do not use decimals. Thanks for your help.
Answer by stanbon(75887)

Given that a-3b/a+2b=3, find the value of a+3b/a-2b. Express your answer as an integer or common or improper fraction reduced to lowest terms. Do not use decimals.
------
(a-3b)/(a+2b) = 3
3a+6b = a-3b
2a + 9b = 0
2a = -9b
a = -4.5b
-------
a+3b = -1.5b
a-2b = -6.5b
------
Then (a+3b)/(a-2b) = -1.5b/-6.5b = 15/65 = 3/13
----
Ans:: (a+3b)/(a-2b) = 3/13
=================================
Cheers,
Stan H.